# Beyond Functional Correctness: Investigating Coding Style Inconsistencies in Large Language Models

### v1

@@ 4. Evaluation
  • RQ1: What are the types of coding style inconsistencies between Code LLMs and human?
@@
    • RQ2.b: What are the inconsistency type numbers present in a single code sample?
@@
  • RQ3: Which coding style is better, model-generated code or the ground truth code?
  • RQ4: Can prompting techniques improve the coding style of Code LLMs?
@@ 4.1. RQ1: Coding Style Inconsistency Identification
To identify the inconsistencies in coding styles of Code LLMs and human programmers, we manually analyze the outputs of the four code LLMs. By comparing these outputs with the ground truth, we summarize the types of coding style inconsistencies.
We conduct open coding (Khandkar, 2009) on the code generated by Code LLMs. Initially, we describe the data collection process, followed by a detailed explanation of the coding protocol.
@@ 4.1.1. Data Collection
Model generation. For each of the 230 Python code generation tasks from CoderEval (Yu et al., 2024), we prompt the four Code LLMs to perform code generation using the same prompting template. For each task, we instruct each model to generate 10 results, resulting in an initial total of 2,300 code samples for each model.
Automatic filtering. To ensure the correctness of the collected code samples, we further filter out code samples that fail to pass any of the associated unit tests for the task, leading to 456, 189, 365, 497 results that pass all tests for CodeLlama-7B, StarCoder2-7B, DeepSeekCoder-1.3B, and DeepSeekCoder-6.7B, respectively. We further merge identical code samples to reduce analysis effort, resulting in 1,159 unique samples. We only annotate 1159 unique code samples to ensure that the annotation results for the same code sample generated by different models are consistent, thereby avoiding the situation where the same code sample generated by different Code LLMs is annotated with different results.
@@
Manual filtering. To ensure the quality of collected code samples, we manually check and filter them based on the following three criteria: 1 Style consistency. We filter out results that exhibit no inconsistency in coding style. For example, Figure 1 shows an example of consistent coding style between the code sample generated by Code LLM and corresponding ground truth of a given task. As a result, 56 code samples are filtered out in this way. 2 Functional correctness. We filter out results that implement the task incorrectly despite passing the unit tests. The functional correctness of the generated result is verified by comparing it to the ground truth and the task descriptions. Previous work has shown that existing benchmarks suffer from test sufficiency issues, meaning that even if a generated result passes all tests, there is still a chance it could be incorrect (Liu et al., 2023b). For example, Figure 2 shows an example of wrong implementation generated by LLMs although passing test cases. As a result, 264 code samples are filtered out in this way. 3 Implementation conciseness. We filter out results that contain extra code that does not contribute to fulfilling the function’s implementation requirements (e.g., two exactly the same loops). As a result, 19 code samples are filtered out in this way.
As a result, we obtain 820 unique code samples for the study, with each code sample corresponding to a task and a ground truth. The numbers of samples that passed test cases are 456, 189, 365, and 497 for CodeLlama-7B, StarCoder2-7B, DeepSeekCoder-1.3B, and DeepSeekCoder-6.7B, respectively. These code samples implement the function correctly but exhibit inconsistencies with the ground truth in coding style, constituting the population for performing open coding.
@@ 4.1.2. Data Annotation
We adopt the definitions and classifications of coding style inconsistencies in previous work (Chen and Abedjan, 2023) as the initialization of our classification and conduct open coding (Khandkar, 2009) on the generated results, e.g., the ordering of the code blocks. Our objective is to refine and expand these definitions and classifications to capture detailed instances of coding style inconsistencies for Code LLMs.
Iterative coding. We analyze the code samples one by one. For each code sample, we compare it with the ground truth line by line to identify the inconsistencies, without knowing which model produced the result. If a code sample and its corresponding ground truth show inconsistency that matches a current definition of inconsistency type, we code the generated result with the specific inconsistency type. If the inconsistency does not fit any existing definitions, we either modify an existing definition or create a new type. When the inconsistency types are updated, all code samples will be re-annotated to ensure consistency. Note that a code sample can be classified under multiple inconsistency types. For example, if a code sample uses a different naming convention (Naming Formatting Inconsistency) and also structures loops differently (Loop Structure Inconsistency), it will be annotated with both inconsistency types.
This iterative coding process aims to capture the nuanced nature of coding style inconsistencies. During the coding process, we also summarize guidelines for each inconsistency type annotation to ensure clarity and consistency in our annotations. These guidelines include specific examples and detailed descriptions to help identify and classify each type of inconsistency accurately. This ensures the annotation consistency and the reproducibility across different coders.
Periodic review and update. After analyzing every 50 code samples, we conduct a review of both the inconsistency type terminology and the coded samples. Based on insights from the review and discussions, we refine the definitions of inconsistency types, merging or removing types as necessary. Following any updates to the terminology, all code samples are re-annotated to maintain consistency and accuracy in the categorization of inconsistencies. This periodic review and update process continues until all code samples have been fully coded, ensuring thorough and reliable identification of coding style inconsistencies. Note that the terminology has remained stable during the last several reviews, indicating a mature and robust classification system. Three of the authors perform the manual filtering and the coding together, resolving disagreements through discussions.
@@ 4.1.3. Taxonomy
Figure 3. Coding Style Inconsistency Terminology.
Figure 3 presents the 24 inconsistency types identified during the open coding, along with their names and definitions. For each inconsistency type, the full annotation results and detailed annotation guidelines are included in our replication package (rep, 2024). We have further categorized the 24 types of inconsistencies into five dimensions based on their main focus:
@@
Figure 4 provides a visual representation of the relationships between the five dimensions and the 24 inconsistency types identified. The inconsistency types are organized into a tree-like structure in the figure, with the dimensions and inconsistency types represented using different shapes, connected by lines. Those inconsistency type sharing the same color indicate they belong to the same dimension. Furthermore, these inconsistencies vary in their scopes of influence, such as identifier, statement, and block, as also depicted in Figure 4. Some inconsistencies may belong to only one or a few identifiers (e.g., Naming Formatting Inconsistency) or a single statement (e.g., Assignment Style Inconsistency), while others may impact an entire block of code (e.g., Loop Structure Inconsistency) or span across multiple blocks (e.g., Code Order Inconsistency). Note that certain inconsistencies could affect both statement and block structures, contingent upon the complexity of the code involved. For instance, in the context of API usage inconsistency, the implementation of the same functionality may vary. It could involve calling different single APIs within a statement, or it might require the coordination of several APIs with specific usage patterns across multiple code blocks.
Compared with the coding style taxonomy of Chen et al. (Chen and Abedjan, 2023), they categorize coding styles into text style and structure style, with four subtypes formatting, naming, ordering of code blocks, and control structures. Our terminology covers all these types and introduces three additional dimensions: semantic, expression/statement, and fault tolerance. We expand upon their framework by introducing 24 fine-grained types compared to 4 types. For instance, we refine their subtype Control Structures into three specific inconsistency types related to: Conditional Structure Inconsistency, Loop Structure Inconsistency, and Control Flow Structure Inconsistency, offering a more detailed classification. Our terminology is backed by comprehensive guidelines derived from actual open coding, providing detailed and actionable classifications.
In summary, our terminology not only complements but also substantially enhances previous research, filling critical gaps and offering a more robust framework for analyzing the inconsistencies in coding style. Note that while our terminology is based on summarizing inconsistencies observed in Python code generated by Code LLMs, it is not limited to Python alone. The concepts and categories can be generalized to other programming languages as needed.
@@ 4.2. RQ2: Coding Style Inconsistency Analysis
We design RQ2 to evaluate the differences between human-written code and CodeLLM-generated code. Specifically, we investigate the coding style differences in three perspectives: (1) Percentages of inconsistent coding styles; (2) Inconsistency numbers present in a single code sample; and (3) Distribution of coding style inconsistency types.
Figure 5. Percentages of Inconsistent Coding Styles. DSC-1.3B and DSC-6.7B are short for DeepSeekCoder-1.3B and DeepSeekCoder-6.7B, respectively.
@@ 4.2.1. Percentages of Inconsistent Coding Styles
Figure 5 shows the percentages of inconsistent coding styles for each Code LLM. The initial number of functionally correct code samples (before deduplication) produced by the four Code LLMs (CodeLlama-7B, StarCoder2-7B, DeepSeekCoder-1.3B, and DeepSeekCoder-6.7B) are 391, 142, 277, and 375, respectively.
From Figure 5, we can find that all code LLMs exhibit coding style inconsistency with human and the inconsistency degree varies: 66.2%, 82.4%, 88.5%, and 89.9% for CodeLlama-7B, StarCoder2-7B, DeepSeekCoder-1.3B, and DeepSeekCoder-6.7B, respectively.
@@ 4.2.2. Inconsistency Numbers Present in a Single Code Sample
Figure 6. Inconsistency Numbers in a Single Code Sample.
For each model, we counted the number of inconsistency types present in each code sample. Then, We counted the frequency of different numbers of inconsistent types in one sample for each model. A line chart was plotted based on the frequency of inconsistency types present in the code samples. From Figure 6, it can be seen that the number of inconsistent types for one code sample ranges between 1 and 9. For each model, the trend of the frequency line chart is roughly the same, with all lines generally showing a decreasing trend. Among them, the code samples of the models all have the highest frequency of having 1 inconsistency type, at 34%, 28%, 38%, and 37% respectively. The lowest frequency is that code samples with 9 inconsistency types, at 1%, 1%, 1%, and 0% respectively.
@@ 4.2.3. Distribution of Coding Style Inconsistency Types
Figure 8. Breakdown Distribution of Coding Style Inconsistency (by Model).
@@
Figure LABEL:fig:distribution:overall illustrates the overall inconsistency distribution in different models. We can observe that the top-4 inconsistency types are API Usage (270.7%), Blank Line (99.2%), Comment Formatting (86.8%), and Data Structure Construction (86.6%), significantly higher than other inconsistency types. Among these top four inconsistency types, API Usage Inconsistency stands out with a significantly higher frequency, even surpassing the combined frequencies of the second and third-ranked types. In contrast, the bottom inconsistency types are: Comment Semantics Inconsistency, Loop Structure Inconsistency, Runtime Validation Inconsistency, Space Inconsistency, Statement Organization Inconsistency, and Input Validation Inconsistency. The low frequencies in these types indicate that Code LLMs and human-written code are relatively consistent in these aspects.
@@
In order to understand the inconsistencies deeper, we conducted a detailed analysis of the top-4 inconsistency types. In our observed code samples and corresponding ground truths, we found that the code samples and corresponding ground truths might call functions from different sources and in varying quantities to achieve similar functionality. Different sources refer to functions that may be defined within the original repository, built-in Python functions, etc. For example, we found that in 6.6% of cases, the ground truth calls functions defined in the original repository while similar functionality is achieved using Python built-in functions, etc., in the code samples generated by models. This may be because the model lacks contextual information about the functions defined in the original repository when generating code. As a result, the large model uses built-in functions or third-party library functions, etc., to achieve similar functionality. For instance, in one task, the ground truth uses a function defined in the original repository, “match_file_by_prefix(prefix, file)”, to check if the prefix of the file name is “prefix”, while the code sample generated by models uses the built-in method in Python “startswith” to achieve similar functionality.
@@
Blank Lines Inconsistency and Comment Format Inconsistency are the second and third most frequent inconsistency types. Among them, the four models show similar frequencies in the category of blank lines. In our observed code samples and corresponding ground truths, we found that, compared to code written by human programmers, the code samples generated by models shows a preference against using blank lines to separate code blocks. The four models generally have a high frequency of Comment Formatting inconsistency, but there are differences among them (StarCoder2-7B has the highest frequency at 35.9%, CodeLlama-7B is second at 23.6%, DeepSeekCoder-1.3B and DeepSeekCoder-6.7B have the lowest frequencies at 14.3% and 13.1%, respectively). The reason for the high frequency of Comment Formatting Inconsistency across the four models is that, in our observed code samples and corresponding ground truths, the code generated by models shows a preference against generating semantically meaningful inline comments compared to the code written by human programmers. One reason for the large frequency difference between StarCoder2-7B and CodeLlama-7B compared to DeepSeekCoder-1.3B and DeepSeekCoder-6.7B is that the comment formatting in the code samples generated by DeepSeekCoder-1.3B and DeepSeekCoder-6.7B is more standard than that in the code samples generated by StarCoder2-7B and CodeLlama-7B. For example, the code samples generated by CodeLlama-7B and StarCoder2-7B may contain commented-out code or TODO comments, while the code samples generated by DeepSeekCoder-1.3B and DeepSeekCoder-6.7B do not. We consider that having commented-out code in code is not good coding practice because these comments are unnecessary information and do not help in understanding the functionality of the code. We also consider including TODO comments is not good coding practice, because high-quality code should be self-explanatory. This means that the code itself should be clear and understandable without the need for additional comments indicating unfinished tasks or future improvements.
Data structure construction inconsistency is a frequently occurring type of inconsistency. The code samples and the corresponding ground truths may show differences in constructing data structures (e.g., list, set). In our observed samples, human programmers tend to prefer using list comprehensions to construct lists, whereas the code samples generated by Code LLMs tends to favor conventional methods for constructing lists.
@@
Figure 9 shows a radar chart of the frequency of inconsistency types for four different models, allowing us to compare the overall frequency distribution of inconsistency types across different models. As shown in Figure 9, the distribution of inconsistency types for DeepSeekCoder-1.3B and DeepSeekCoder-6.7B is relatively similar compared to CodeLlama-7B and StarCoder2-7B. For example, in the Inline Code Usage inconsistency type, the frequency for DeepSeekCoder-1.3B and DeepSeekCoder-6.7B is higher compared to CodeLlama-7B and StarCoder2-7B. In our observed samples, both DeepSeekCoder-1.3B and DeepSeekCoder-6.7B tend to include more intermediate variables in their code compared to the ground truths. Therefore, we can conclude that the base model significantly influences the coding style. The training data and method have a more noticeable impact on the coding style of the model compared to the parameters.
@@
Figure 9 presents a radar chart that summarizes coding style inconsistencies by grouping them into five broader dimensions, i.e., formatting, semantic, expression/statement, control flow, and fault tolerance. To calculate the frequency for each dimension, we sum the instances of inconsistency types belonging to that dimension and divide it by the total number of valid code samples.
From Figure 9, we have the following observations:
@@
  • It is evident that the coding styles of different Code LLMs are similar in dimension granularity. This is indicated by the almost overlapping shapes on the radar chart, highlighting that these models share a similar distribution of inconsistency types by dimension.
@@
  • The dimensions, ranked by average frequency of inconsistencies, are as follows: statement/expression (73.7%), formatting (49.9%), fault tolerance (24.2%), control flow (17.4%), and semantic(6.3%). The high ranking of statement/expression inconsistency is primarily due to the significantly high frequency of API Usage Inconsistency within this dimension.

### v2

@@ 4. Evaluation
  • RQ1: What are the types of coding style inconsistencies between the selected LLMs and human?
@@
    • RQ2.b: What are the inconsistency type numbers present between a single code sample and the corresponding ground truth?
@@
  • RQ3: How do model-generated code and human-written code compare in three aspects: readability, conciseness, and robustness?
  • RQ4: Can prompting techniques improve the coding style of LLMs?
@@ 4.1. RQ1: Coding Style Inconsistency Identification
To identify the inconsistencies in coding styles between LLMs and human programmers, we manually analyze the generated results of the five LLMs. It’s worth noting that human-written code does not always equate to good coding style. Therefore, we do not use human-written code as a standard to assess the quality of coding style in LLM-generated code. Instead, we focus solely on analyzing the coding style inconsistencies between human-written code and LLM-generated code. By comparing these generated results with the ground truths, we summarize the types of coding style inconsistencies. We conduct open coding (Khandkar, 2009) on the code generated by LLMs. Initially, we describe the data collection process, followed by a detailed explanation of the coding protocol.
@@ 4.1.1. Data Collection
Model generation. For each of the 230 Python code generation tasks from CoderEval (Yu et al., 2024), we prompt the five LLMs to perform code generation using the same prompting template. For each task, we instruct each model to generate 10 results, resulting in an initial total of 2,300 code samples for each model.
Automatic filtering. To ensure the correctness of the collected code samples, we further filter out code samples that fail to pass any of the associated unit tests for the task, leading to 456, 189, 365, 497 and 570 results that pass all tests for five LLMs, respectively. We further merge identical code samples to reduce analysis effort, resulting in 1,557 unique code samples. We only annotate these unique code samples to ensure that the annotation results for the same code sample generated by different models are consistent, thereby avoiding the situation where the same code sample generated by different LLMs is annotated with different results.
@@
Manual filtering. To ensure the quality of collected code samples, we manually check and filter them based on three criteria: 1 Style consistency. We filter out results that exhibit no inconsistency in coding style. If the naming conventions, commenting style, code structure, and other aspects of the two code samples are consistent, we conclude that there is no significant difference in code style between them. For example, Figure 2 shows an example of consistent coding style between the code sample generated by Code LLM and corresponding ground truth. As a result, 73 code samples are filtered out in this way. 2 Functional correctness. We filter out results that implement the task incorrectly despite passing the unit tests. Previous work has shown that existing benchmarks suffer from test sufficiency issues, meaning that even if a generated result passes all tests, there is still a chance it could be incorrect (Liu et al., 2023a). For example, Figure 2 shows an example of wrong implementation although passing test cases. As a result, 286 code samples are filtered out in this way. 3 Implementation conciseness. We filter out results that contain extra code that does not contribute to fulfilling the function’s implementation requirements (e.g., two exactly the same loops). As a result, 19 code samples are filtered out in this way. Finally, we obtain 1,179 unique code samples for the study.
@@ 4.1.2. Data Annotation
We adopt classifications of coding style inconsistencies in previous work (Chen and Abedjan, 2023) as the initialization of our classification and conduct open coding (Khandkar, 2009) on the code samples generated by LLMs. Our objective is to refine and expand these classifications to capture detailed instances of coding style inconsistencies. In our annotation process, three annotators were involved, each with over five years of Python programming experience and a solid understanding of the field of code style.
Iterative coding. The three annotators independently analyze the code samples one by one. For each code sample, the annotators independently compare it with the ground truth line by line to identify inconsistencies, without knowing which model produced the result. If a code sample and its corresponding ground truth show inconsistency that matches a current definition of inconsistency type, the annotators code the generated result with the specific inconsistency type. If the inconsistency does not fit any existing definitions, the annotators either modify an existing definition or create a new type. When the inconsistency types are updated, all code samples will be re-annotated to ensure consistency. Note that a code sample can be classified under multiple inconsistency types. For example, if a code sample follows a different naming convention (Naming Formatting Inconsistency) and also structures loops differently (Loop Structure Inconsistency), it will be annotated with both inconsistency types.
This iterative coding process aims to capture the nuanced nature of coding style inconsistencies. During the coding process, the annotators also summarize guidelines for each inconsistency type annotation to ensure clarity and consistency in our annotations. These guidelines include specific examples and detailed descriptions to help identify and classify each type of inconsistency accurately. This ensures the annotation consistency and the reproducibility across different coders.
Periodic review and update. After analyzing every 50 code samples, the three annotators conduct a review of both the taxonomy and the coded samples. Based on insights from the review and discussions, we refine the definitions of inconsistency types, merging or removing types as necessary. Following any updates to the taxonomy, all code samples are re-annotated to maintain consistency and accuracy in the categorization of inconsistencies. This periodic review and update process continues until all code samples have been fully coded, ensuring thorough and reliable identification of coding style inconsistencies. Note that the taxonomy has remained stable during the last several reviews, indicating a mature and robust classification system. We calculated that the average disagreement rate among the three annotators was 37.6% after labeling every 50 samples. However, these disagreements were resolved through discussions among the annotators.
@@ 4.1.3. Taxonomy
Table 1 presents the 24 inconsistency types identified during the open coding, along with their names and definitions. For each inconsistency type, the full annotation results and detailed annotation guidelines are included in our replication package (rep, 2024).
Taxonomy Analysis. We have further categorized the 24 types of inconsistencies into five dimensions based on their main focus:
@@
It is worth noting that we have categorized 24 inconsistency types into 5 dimensions, with the scope of code style inconsistencies contained in each dimension solely determined by the inconsistency types assigned to that dimension. Since there is no obvious overlap between the inconsistency types, there is also no clear overlap between the 5 dimensions.
Figure 3 provides a visual representation of the relationships between the five dimensions and the 24 inconsistency types identified. The inconsistency types are organized into a tree-like structure in the figure, with the dimensions and inconsistency types represented using different shapes, connected by lines. Furthermore, these inconsistencies vary in their scopes of influence, such as identifier, statement, and block, as also depicted in Figure 3. Some inconsistencies may belong to only one or a few identifiers (e.g., Naming Formatting Inconsistency) or a single statement (e.g., Assignment Style Inconsistency), while others may impact an entire block of code (e.g., Loop Structure Inconsistency) or span across multiple blocks (e.g., Code Order Inconsistency).
Table 1. Coding Style Inconsistency taxonomy.
ID Inconsistency Type Definition
1 Naming Format Inconsistency Inconsistencies in the formatting of identifiers (e.g., variable names, function names, or parameter names), such as using camelCase (e.g., authorName) versus snake_case (e.g., author_name).
2 Space Inconsistency Inconsistencies in the use of space(e.g., whitespace and indentation) around various syntactical elements, e.g., operators, colons, comments, and brackets.
3 Blank Line Inconsistency Inconsistencies in the use of blank lines. For example, one style includes blank lines to separate code blocks, while the other omits them.
4 Inline Code Usage Inconsistency Inconsistencies in the usage of inline code constructs. It encompasses cases where one approach employs inline expressions or functions while the other does not.
5 Comment Format Inconsistency Inconsistencies in the formatting of comments within code. It includes variations in interline comments, inline comments, commented-out code, and trailing comments.
6 Statement Organization Inconsistency Inconsistency in the organization style of statements, exemplified by completing expressions or statements in a single line in contrast to breaking them into multiple shorter lines.
7 Naming Semantics Inconsistency Inconsistencies in the semantic meaning of identifiers, such as using generic single-letter identifiers (e.g., i, l, d) versus meaningful, descriptive words (e.g., index, length, day).
8 Comment Semantics Inconsistency Inconsistencies in the semantic aspects of comments within code, such as variations in the level of detail or semantic differences, or with TODO comment, useless comments.
9 Assignment Inconsistency Inconsistencies in the style of variable assignment, e.g., tuple unpacking assignment, chained assignment, separate assignment. Examples include using augmented assignment versus standard assignment (e.g., ‘x += 1‘ vs. ‘x = x + 1‘).
10 Conditional Syntax Inconsistency Inconsistencies in the syntax used for conditional statements within code. It covers scenarios where one method involves conditional statements while the other employs conditional expressions or return statements with equivalent functionality.
11 Conditional Expression Inconsistency Inconsistencies in the way conditional expressions are written, despite having similar functionalities. For example, one style might use if len(a) ¿ 1 while another uses if len(a) ¿= 2.
12 Data Structure Construction Inconsistency Inconsistencies in the methods used to construct data structures such as lists, dictionaries, sets, tuples, strings, and iterators. For example, using different syntaxes or functions to create these data structures.
13 API Preference Inconsistency Inconsistencies in how APIs are used to achieve similar functionality. It includes variations such as calling different functions or methods defined in the repository, using built-in functions, or re-implementing the functionality without calling existing functions.
14 Advanced Syntax Usage Inconsistency Inconsistencies in the use of advanced syntax features, such as lambda expressions
15 Code Ordering Inconsistency Inconsistencies in the order of semantically similar code blocks, such as import statements, assignments, loops, and other logical sections of code.
16 Loop Structure Inconsistency Inconsistencies in loop structures within code. It covers scenarios where one approach employs a for loop while the other uses a while loop, or where one loop contains only a basic loop structure while the other includes additional control flow statements such as if-break, for-else structure, and while-else structure
17 Conditional Structure Inconsistency Inconsistencies in the structure and design of conditional statements within methods. It includes variances such as the use of multiple conditional statements versus a single statement with equivalent meaning, differences in the structures of multiple conditional statements while preserving the same semantics, and disparities in the inclusion of return statements alongside conditional statements.
18 Control Flow Structure Inconsistency Inconsistencies in the use of control flow structures, such as using if-else versus try-except statements during the code execution (e.g., input and runtime validation).
19 Input Validation Presence Inconsistency Inconsistencies in whether input checking with conditionals is performed.
20 Runtime Validation Presence Inconsistency Inconsistencies in whether runtime validation with conditionals is performed, ensuring data integrity during code execution..
21 Exception Handling Presence Inconsistency Inconsistencies in whether exceptions are handled, e.g., using try-except blocks, to manage errors that occur during execution.
22 Input Validation Style Inconsistency Inconsistencies in the style of input validation with conditionals (ensuring input data is checked before processing), such as whether exceptions are thrown, the types of exceptions used, and the use of logging.
23 Runtime Validation Style Inconsistency Inconsistencies in the style of runtime validation with conditionals during code execution, such as whether exceptions are thrown, the types of exceptions used, and the use of logging.
24 Exception Handling Style Inconsistency Inconsistencies in the style of exceptions that occur during execution are handled, such as whether exceptions are thrown, the types of exceptions used, and the use of logging, the use of try-else block.
@@
Note that certain inconsistencies could affect both statement and block structures, contingent upon the complexity of the code involved. For instance, in the context of API preference inconsistency, the implementation of the same functionality may vary. It could involve calling different single APIs within a statement, or it might require the coordination of several APIs with specific usage patterns across multiple code blocks.
Taxonomy Comparison. Compared with the coding style taxonomy of Chen et al. (Chen and Abedjan, 2023), they categorize coding styles into text style and structure style, with four subtypes formatting, naming, ordering of code blocks, and control structures. Our taxonomy covers all these types and introduces three additional dimensions: semantic, expression/statement, and fault tolerance. We expand upon their framework by introducing 24 fine-grained types compared to 4 types. For instance, we refine their subtype Control Structures into three specific inconsistency types related to: Conditional Structure Inconsistency, Loop Structure Inconsistency, and Control Flow Structure Inconsistency, offering a more detailed classification. Our taxonomy is backed by comprehensive guidelines derived from actual open coding, providing detailed and actionable classifications.
Taxonomy Generalizability. One concern is that, since each task in CoderEval has only one ground truth, this may reflect the coding style of a single programmer, potentially limiting the generalizability of our taxonomy. To this end, we randomly selecting 42 Python tasks from the 230 available in CoderEval and two annotators were assigned to independently rewrite the ground truths for these 42 tasks, resulting in a total of 84 ground truths. Another annotator was responsible for verifying the correctness of these ground truths. We then identified the code samples corresponding to these 42 tasks from a set of 1,179 unique code samples and annotated the coding style differences between the rewritten ground truths and the code samples following the outlined annotation process. Both the rewritten ground truths and the annotation results are available in our replication package (rep, 2024). After the annotation process, the overall taxonomy remained unchanged.
In summary, our taxonomy not only complements but also substantially enhances previous research, filling critical gaps and offering a more robust framework for analyzing the inconsistencies in coding style. Note that while our taxonomy is based on summarizing inconsistencies observed in Python code generated by LLMs, it is not limited to Python alone. The concepts and categories can be generalized to other programming languages as needed.
@@ 4.2. RQ2: Coding Style Inconsistency Analysis
We design RQ2 to evaluate the differences between human-written code and LLM-generated code. Specifically, we investigate the coding style differences in three perspectives: (1) Percentages of inconsistent coding styles; (2) Inconsistency numbers present in a single code sample; and (3) Distribution of coding style inconsistency types.
Figure 4. Percentages of Inconsistent Coding Styles.
Figure 5. Inconsistency Numbers in a Single Code Sample.
@@ 4.2.1. Percentages of Inconsistent Coding Styles
We counted the number of code samples generated by the five models (CodeLlama-7B, StarCoder2-7B, DeepSeekCoder-1.3B, DeepSeekCoder-6.7B, and GPT-4) that passed unit tests, were manually verified for functional correctness, and did not contain extra code. The results were 391, 142, 277, 375, and 537, respectively. Furthermore, we calculated the proportion of these code samples that either exhibit or do not exhibit coding style inconsistencies compared to human-written code. The statistical results are shown in Figure 5. From Figure 5, we can find that all LLMs exhibit coding style inconsistency with human-written code and the degree varies: 66.2%, 82.4%, 88.5%, 89.9% and 90.1% for the five models, respectively.
@@ 4.2.2. Inconsistency Numbers Present in a Single Code Sample
For each model, we counted the number of inconsistency types present in each code sample (considering only the code samples that exhibit coding style differences with the corresponding ground truths). Then, We counted the frequency of different numbers of inconsistent types in one sample for each model. A line chart was plotted based on the frequency of inconsistency types present in the code samples. From Figure 5, it can be seen that the number of inconsistent types for one code sample ranges between 1 and 8. For each model, the trend of the frequency line chart is roughly the same, with all lines generally showing a decreasing trend. The code samples of CodeLlama-7B, StarCoder2-7B, DeepSeekCoder-1.3B, and DeepSeekCoder-6.7B show the highest frequency of having one inconsistency type, at 39.8%, 32.5%, 35.9%, and 37.8%, respectively. The code samples of GPT-4 exhibit the highest frequency of two inconsistency types, at 31.8%. The lowest frequency is that code samples with 8 inconsistency types, at 0.0%, 0.0%, 0.8%, 0.3% and 0.0%, respectively.
@@ 4.2.3. Distribution of Coding Style Inconsistency Types
Figure 7. Breakdown Distribution of Coding Style Inconsistency (by Model). The “Semantics” at the top refers to “Naming Semantics”, and the “Presence”at the bottom refers to “Input Validation Presence”.
@@
Figure 8 is a stacked plot showing the frequencies of a certain inconsistency type in code samples generated by five models. For example, the frequencies of API Preference inconsistency in the code samples generated by CodeLlama-7B, StarCoder2-7B, DeepSeekCoder-1.3B, DeepSeekCoder-6.7B, and GPT-4 are 70.7%, 63.3%, 72.7%, 64.1%, and 54.4%, respectively, summing up to 325.0%. We can observe that the top-4 inconsistency types are API Preference Inconsistency (325.0%), Comment Formatting Inconsistency (126.3%), Blank Line Inconsistency (126.1%) and Data Structure Construction Inconsistency (97.4%). Among these top four inconsistency types, API Preference Inconsistency stands out with a significantly higher frequency, even surpassing the combined frequencies of the second and third-ranked types. In contrast, the bottom inconsistency types are: Runtime Validation Style (9.7%), Loop Structure (7.4%), Comment Semantic (2.5%). The low frequencies of these types indicate that LLMs and human-written code are relatively consistent in these aspects.
@@
In order to understand the inconsistencies deeper, we conducted a detailed analysis of the top-4 inconsistency types. In our observed code samples and the ground truths, we found that the code samples and the ground truths might call functions from different sources to achieve similar functionality. Different sources refer to functions that may be defined within the original repository, built-in Python functions, etc. For example, we found that in 6.6% of cases, the ground truth calls functions defined in the original repository while similar functionality is achieved using Python built-in functions, etc., in the code samples generated by models. This may be because the model lacks contextual information about the functions defined in the original repository when generating code. As a result, the large model uses built-in functions or third-party library functions, etc., to achieve similar functionality. For instance, in one task, the ground truth uses a function defined in the original repository, “match_file_by_prefix(prefix, file)”, to check if the prefix of the file name is “prefix”, while the code sample generated by models uses the built-in method in Python “startswith” to achieve similar functionality.
@@
Comment Format Inconsistency and Blank Lines Inconsistency are the second and third most frequent inconsistency types. In our analysis, we found that model-generated code samples tend to avoid using blank lines to separate code blocks compared to human-written code. The five models generally have a high frequency of Comment Formatting inconsistency, but there are differences among them (The highest frequency is observed for GPT-4 at 39.5%. StarCoder2-7B follows with 35.9%, and CodeLlama-7B ranks third at 23.6%. DeepSeekCoder-1.3B and DeepSeekCoder-6.7B have the lowest frequencies, recorded at 14.3% and 13.1%, respectively.). Among these models, GPT-4 has the highest frequency, partly because it includes more meaningful block and inline comments compared to human-written code. The other four models are less inclined to add such comments. Notably, code samples from CodeLlama-7B, StarCoder2-7B, and DeepSeekCoder-1.3B even do not include any examples containing inline comments. The relatively high frequency of StarCoder2-7B and CodeLlama-7B can be attributed to their less standardized comment formatting compared to the other models. For example, the code samples generated by CodeLlama-7B and StarCoder2-7B may contain commented-out code or TODO comments, while the code samples generated by DeepSeekCoder-1.3B, DeepSeekCoder-6.7B and GPT-4 do not. We consider that having commented-out code in code is not good coding practice because these comments are unnecessary information and do not help in understanding the functionality of the code. We believe that including TODO comments in code generated by large models is not good coding practice. This is because we expect large models to produce complete code based on requirements, rather than including comments indicating unfinished tasks or future improvements. Data structure construction inconsistency is a frequently occurring type of inconsistency. The code samples and the corresponding ground truths may show differences in constructing data structures (e.g., list, set). In our observed samples, human programmers tend to prefer using list comprehensions to construct lists, whereas the code samples generated by LLMs tends to favor conventional methods for constructing lists.
@@
Figure 8 shows a radar chart of the frequency of inconsistency types for five different models, allowing us to compare the overall frequency distribution of inconsistency types across these models. As shown in Figure 8, the distribution of inconsistency types for DeepSeekCoder-1.3B and DeepSeekCoder-6.7B is relatively similar compared to the other models in terms of inconsistency types such as comment format, conditional expression, and naming format, etc. For example, in the Inline Code Usage inconsistency type, the frequency for DeepSeekCoder-1.3B and DeepSeekCoder-6.7B is very close, lower than the frequency for GPT-4, but higher than that for CodeLlama-7B and StarCoder2-7B. The frequencies of DeepSeekCoder-1.3B, DeepSeekCoder-6.7B, and GPT-4 are higher because they tend to include more intermediate variables in the code compared to the ground truths. Therefore, we can conclude that the base model significantly influences the coding style. The training data and method have a more noticeable impact on the coding style of the model compared to the parameters.
@@
Figure 8 presents a radar chart that summarizes coding style inconsistencies by grouping them into five broader dimensions, i.e., formatting, semantic, expression/statement, control flow, and fault tolerance. To calculate the frequency for each dimension, we sum the instances of inconsistency types belonging to that dimension and divide it by the total number of valid code samples. From Figure 8, we have the following observations:
@@
  • It is evident that the coding styles of different LLMs are similar in dimension granularity. This is indicated by the almost overlapping shapes on the radar chart, highlighting that these models share a similar distribution of inconsistency types by dimension.
@@
  • The dimensions, ranked by average frequency of inconsistencies, are as follows: statement/expression (73.1%), formatting (52.4%), fault tolerance (24.3%), control flow (18.0%), and semantic(7.5%). The high ranking of statement/expression inconsistency is primarily due to the significantly high frequency of API Preference Inconsistency within this dimension.